# ReAcTree: Hierarchical LLM Agent Trees with Control Flow for Long-Horizon Task Planning

### v1

@@ Abstract.
Recent advancements in large language models (LLMs) have enabled significant progress in decision-making and task planning for embodied autonomous agents. However, most existing methods still struggle with complex, long-horizon tasks because they rely on a monolithic trajectory that entangles all past decisions and observations, attempting to solve the entire task in a single unified process. To address this limitation, we propose ReAcTree, a hierarchical task-planning method that decomposes a complex goal into more manageable subgoals within a dynamically constructed agent tree. Each subgoal is handled by an LLM agent node capable of reasoning, acting, and further expanding the tree, while control flow nodes coordinate the execution strategies of agent nodes. In addition, we integrate two complementary memory systems: each agent node retrieves goal-specific, subgoal-level examples from episodic memory and shares environment-specific observations through working memory. Experiments on the WAH-NL and ALFRED datasets demonstrate that ReAcTree consistently outperforms strong task-planning baselines such as ReAct across diverse LLMs. Notably, on WAH-NL, ReAcTree achieves a 61% goal success rate with Qwen 2.5 72B, nearly doubling ReAct’s 31%.
@@
Task planning, large language models, decision-making, hierarchical planning, behavior trees, embodied agents
@@
1. Introduction
Large language models (LLMs) have recently exhibited remarkable reasoning capabilities by generating intermediate reasoning steps Wei et al. (2022); Kojima et al. (2022). These advances open up new possibilities for decision-making and task planning in embodied autonomous agents, with the long-term objective of enabling them to fulfill high-level natural language commands autonomously. As early progress toward this goal, several works Huang et al. (2022a); Ahn et al. (2022); Song et al. (2023); Liang et al. (2023) have demonstrated that LLMs can leverage their pre-trained world knowledge to generate mid-level action sequences without additional training, primarily through prompting rather than handcrafted heuristics or learned policies Eysenbach et al. (2019); Xu et al. (2019).
The decision-making and planning capabilities of LLMs can be further improved by incorporating observations or feedback from external environments. Recent approaches explore adapting the next action Yao et al. (2023); Huang et al. (2022b); Zeng et al. (2023), adjusting the entire plan Shinn et al. (2024), or refining both the plan and the next action Wang et al. (2023a); Sun et al. (2023). However, these methods frequently use a single, monolithic trajectory that entangles all past decisions and observations from multiple subgoals, which increases the risk of hallucination and logical failures Liu et al. (2024b). Another line of research investigates multiple reasoning paths Wang et al. (2023b); Yao et al. (2024); Zhou et al. (2024) for general reasoning tasks. Such methods typically assume the capability to simulate multiple paths and revert to previous states Zhao et al. (2024); Zhuang et al. (2024); Hu et al. (2024), which is rarely feasible in real-world scenarios due to irreversible actions (e.g., slicing an apple) and partial observability.
@@ 1. Introduction
To tackle these challenges, we introduce ReAcTree, a novel framework that dynamically constructs an agent tree in the subgoal space rather than an action tree over primitive actions. Each LLM-powered agent node is responsible for a subgoal and extends the ReAct paradigm Yao et al. (2023): it can reason, act, or expand the tree by proposing new subgoals with an associated control flow when a task is too complex. Control flow nodes, inspired by Behavior Trees Colledanchise et al. (2018), coordinate these agents by sequencing, fallback, or parallel execution. Conceptually, ReAcTree extends the Least-to-Most prompting strategy Zhou et al. (2023) from static reasoning to dynamic, agentic planning. By decomposing a problem into semantically isolated subgoals and providing each with its own focused context, ReAcTree prevents the propagation of reasoning errors and makes long-horizon tasks more tractable for LLMs. The hierarchical structure of ReAcTree is illustrated in Figure 1.
Figure 1. An illustrative example of how ReAcTree generates an agent tree for the natural language instruction: Please bring one pudding and one juice to the coffee table. The left side shows the hierarchical structure with agent nodes (circles) and control flow nodes (squares); the number inside each circle denotes the execution order, and the attached text box specifies the corresponding subgoal. Control flow nodes are labeled by their types ($\rightarrow$ for sequence, ? for fallback, and $\Rightarrow$ for parallel). The right side highlights the decision-making trajectory of agent node 3, including observation, reasoning, acting, subgoal expansion, and the use of episodic and working memory.
@@
To enhance the in-context learning and coordination capabilities of the proposed agent architecture, we incorporate two complementary memory systems. First, episodic memory assists each agent node in effective in-context learning by providing subgoal-level examples from past experiences that are semantically similar to the current subgoal. It can be bootstrapped from a small set of manually collected trajectories and gradually expanded through successful runs. Second, working memory functions as a shared blackboard, allowing agent nodes to exchange critical observations (e.g., the locations of movable objects). This collective situational awareness reduces redundant searches and mitigates hallucinations. Together, these memory systems enable ReAcTree to learn from the past and coordinate in the present, strengthening the robustness of agentic decision-making under partial observability.
@@
We evaluate ReAcTree by extending LoTa-Bench Choi et al. (2024), a benchmark providing two household task environments (WAH-NL Puig et al. (2021); Choi et al. (2024) with VirtualHome Puig et al. (2018) and ALFRED Shridhar et al. (2020) with AI2THOR Kolve et al. (2017)), to a more challenging partially observable setting. This setup reflects real-world constraints and tests agentic decision-making under uncertainty. Across both environments, ReAcTree consistently outperforms strong baselines across various LLMs. On WAH-NL, it achieves a 61% goal success rate with Qwen 2.5 72B, nearly doubling ReAct’s 31% under the same model size. Notably, even with a much smaller 7B model, it reaches 37%, showing that ReAcTree maintains strong performance despite reduced model capacity. The contributions of its core components are further substantiated by ablation studies on memory systems and control flow types, as well as in-depth analyses of computational cost and failure modes.
@@
In summary, this paper makes the following contributions: (1) ReAcTree, a hierarchical planning framework for long-horizon tasks that dynamically constructs an LLM agent tree, where each agent node independently solves a subgoal while control flow nodes coordinate their execution into a coherent plan; (2) two memory systems: episodic memory that enhances in-context learning ability of agent nodes, and working memory that facilitates information sharing across agent nodes to support collective situational awareness; and (3) extensive experiments on LoTa-Bench under partially observable settings, demonstrating the effectiveness of ReAcTree across strong baselines and providing detailed analysis through ablations, computational cost analysis, and failure case studies. The code is available https://github.com/Choi-JaeWoo/ReAcTree.
@@ 2. Related Work
LLM-based Embodied Agents. LLMs such as GPT-3, GPT-4, PaLM, LLaMA, and Qwen Brown et al. (2020); Achiam et al. (2023); Chowdhery et al. (2023); Dubey et al. (2024); Yang et al. (2024) have demonstrated impressive few-shot in-context learning capabilities, advancing the state of the art in NLP tasks. Prompting techniques Wei et al. (2022); Kojima et al. (2022); Zhou et al. (2023); Gao et al. (2023) further enhance reasoning by encouraging articulation of intermediate steps. Building on these advancements, research has increasingly explored text-based applications to embodied decision-making. Early studies Huang et al. (2022a); Ahn et al. (2022); Song et al. (2023) showed that LLMs could generate mid-level action sequences without additional training, while later studies introduced code-based plan generation Liang et al. (2023); Singh et al. (2023), the extension of classical agent architectures Ichida et al. (2024), and the integration of environmental feedback or tools Huang et al. (2022b); Zeng et al. (2023); Shen et al. (2024); Lu et al. (2024); Chen et al. (2023). ReAct Yao et al. (2023) enhanced planning by prompting explicit intermediate reasoning, while Reflexion Shinn et al. (2024) applied iterative self-refinement Madaan et al. (2024); Welleck et al. (2023); Paul et al. (2024); Chen et al. (2024b), offering additional flexibility for long-horizon tasks.
@@
Hierarchical Task Planning with LLMs. To tackle more complex, long-horizon tasks, researchers have introduced hierarchical frameworks that decompose goals into manageable planning levels. AdaPlanner Sun et al. (2023) and DEPS Wang et al. (2023a) adopt bi-level hierarchies, refining both an overall plan and next-step decisions with environmental feedback. Other methods combine classical task-and-motion planning with learned low-level controllers guided by LLM reasoning Wong et al. (2023); Liu et al. (2024a); Ding et al. (2023); Shukla et al. (2024); Wang et al. (2025). Some works leverage behavior trees Wang et al. (2024); Chen et al. (2024a); Ao et al. (2025) to organize high-level actions in specialized domains such as human-robot cooking or robotic assembly. However, they typically rely on predefined structures or domain-specific routines. In contrast, our approach, ReAcTree, enables dynamic task decomposition, expanding subgoals in complex environments without being tied to a specific domain.
@@
Tree Search-Based Planning with LLMs. A prominent line of research explores multiple reasoning paths to evaluate diverse hypotheses before committing to a final solution. Such approaches Wang et al. (2023b); Yao et al. (2024); Zhou et al. (2024); Xie et al. (2024); Besta et al. (2024); Hao et al. (2023) show that systematically branching reasoning steps significantly improves performance in reasoning tasks. Several studies extend this idea to agentic planning. For instance, LLM-MCTS Zhao et al. (2024) and ToolChain* Zhuang et al. (2024) search action trees using Monte Carlo Tree Search and A* search, respectively. Tree-Planner Hu et al. (2024) independently samples planning paths and merges them into an action tree, then makes decisions based on grounded observations at each node. However, these methods typically assume reversible actions (i.e., rolling back to prior states) in simulators, which limits their applicability in real-world settings. In contrast, ReAcTree constructs a dynamically expanding LLM agent tree that decomposes complex goals into subgoals, delegates them to agent nodes, and coordinates their execution via behavior tree-inspired control flow, replacing search-based exploration with agent coordination.
@@ 3. Preliminaries
Problem Formulation. We consider the agentic planning problem as a sequential decision-making problem aimed at achieving a goal $g$ expressed in natural language. At each time step $t$, the agent has access to the context $c_{t}=(o_{1},a_{1},o_{2},a_{2},\cdots,a_{t-1},o_{t})$, where $o_{i}$ and $a_{i}$ represent the observation and action at time step $i$, respectively. The objective of the agent is to generate the next appropriate action $a_{t}$ based on the context $c_{t}$, with the aim of eventually achieving the goal $g$.
ReAct (Yao et al., 2023). ReAct addresses this problem by interleaving reasoning and action execution using a pre-trained LLM $p_{LLM}$. The action policy is defined as $a_{t}\sim p_{LLM}(\cdot|P,g,c_{t})$, where $P=(P_{sys},P_{ic})$ is the initial prompt, composed of a system prompt $P_{sys}$ and in-context examples $P_{ic}$. The key idea is to use the augmented action space, $\hat{\mathcal{A}}_{t}=\mathcal{A}_{t}\cup\mathcal{L}$, where $\mathcal{A}_{t}$ is the set of executable skills available at time $t$, and $\mathcal{L}$ is the language space representing reasoning steps or thoughts. If $a_{t}\in\mathcal{A}_{t}$, the agent executes the action and obtains a text observation from the environment. If $a_{t}\in\mathcal{L}$, it is called a thought or reasoning trace, which aids in the logical inference of the LLM. In this case, the agent does not receive a new observation from the environment, i.e., $o_{t+1}=\phi$.
@@ 4. ReAcTree
ReAcTree is a hierarchical task-planning framework for agentic decision-making in complex environments. It dynamically constructs a tree of agent nodes and control flow nodes, as illustrated in Figure 2. Agent nodes operate as LLM-based task planners that can reason, act, and expand the tree with new subgoals and appropriate control flow. Control flow nodes, inspired by behavior trees (Colledanchise et al., 2018), determine how child agent nodes are executed and how their outcomes are propagated upward in the tree. This design offers two key advantages: (1) isolating subgoals reduces hallucination and logical errors in long trajectories by keeping each agent node focused on its local context and goal, which also enables targeted in-context example selection, and (2) control flow nodes ensure robust and interpretable execution logic for reliable planning.
@@ 4.1. ReAcTree Algorithm
Control Flow Nodes. Each control flow node coordinates the execution of its child agent nodes according to behavior tree principles (Colledanchise et al., 2018). ReAcTree supports three types of control flow nodes: sequence ($\rightarrow$), fallback (??), and parallel ($\Rightarrow$). The sequence node executes its children sequentially, returning success only if all of them succeed and failing immediately if any child fails. The fallback node also executes children in order but returns success as soon as the first child succeeds and fails only if all children fail. Lastly, the parallel node executes all children sequentially without early termination and aggregates outcomes according to a predefined policy; we adopt majority voting, which is useful for subgoals that can be attempted independently (e.g., placing multiple objects in different locations). After execution, a control flow node reports its overall success or failure to its parent node. The complete pseudocode is provided in Appendix A.
@@ 4.2. Memory Systems
To support ReAcTree’s hierarchical planning and decision-making, we introduce two complementary memory systems: episodic memory and working memory. Episodic memory enhances the in-context learning capability of agent nodes by storing and retrieving past subgoal-level experiences. In contrast, working memory facilitates information sharing across nodes by recording environment-specific observations, such as the latest location of movable objects, during task execution.
@@
Episodic Memory. In ReAcTree, each agent node $e$ is responsible for a particular subgoal $g^{e}$, and its entire decision-making trajectory forms a subgoal-level experience. Episodic memory $M_{ep}$ stores experiences from agent nodes involved in task executions that ultimately succeed, even if the individual node did not complete its subgoal.
@@
By storing experiences at the subgoal level, episodic memory maintains shorter, goal-directed trajectories than monolithic planners like ReAct. For instance, while ReAct stores a single long trajectory for Bring pudding and juice to the coffee table, ReAcTree instead stores focused trajectories for each subgoal, such as find and pick up pudding and find and pick up juice. This finer granularity enables more targeted in-context retrieval and improves decision accuracy at each agent node.
@@
Formally, each experience is recorded as a tuple $(t^{e},v^{e},s^{e})$, where: (1) $t^{e}=(g^{e},\,o^{e}_{1},\,a^{e}_{1},\,\dots,\,o^{e}_{T},\,a^{e}_{T})$ denotes the full text trajectory, recording all observations $o^{e}_{t}$ and actions $a^{e}_{t}$ during the node’s lifetime; (2) $v^{e}=f_{\text{sen}}(g^{e})$ is the sentence embedding of the node’s subgoal, computed using a pretrained encoder $f_{\text{sen}}$ such as Sentence-BERT (Reimers and Gurevych, 2019); and (3) $s^{e}\in\{\textit{success},\textit{failure},\textit{expand}\}$ is the termination state of the agent node.
The collected episodic memory $M_{ep}$ provides in-context examples for agent node inference. Before decision-making begins at inference time, an agent node retrieves in-context examples from $M_{ep}$ by comparing its current subgoal $g^{n}$ to the stored goals using cosine similarity. Specifically, the agent embeds its goal $g^{n}$ as $v^{n}=f_{\text{sen}}(g^{n})$ and computes the similarity with each $v^{e}$ in $M_{ep}$:
$\text{sim}(v^{n},\,v^{e})=\frac{v^{n}\cdot v^{e}}{\|\,v^{n}\,\|\;\|\,v^{e}\,\|}.$ (2)
@@
The agent retrieves the top $k$ most similar experiences (with $k$ determined by a token limit) to assemble as in-context examples. If multiple experiences have the same similarity score, ReAcTree samples uniformly across termination states {success, failure, expand} to promote diversity. In practice, episodic memory can be bootstrapped from a few manually collected trajectories and gradually augmented with additional successful executions.
@@
Working Memory. Working memory captures environment-specific information within a single ReAcTree run and is shared across all agent nodes to store and recall key observations. In this work, we focus on tracking the locations of movable objects to reduce redundant interactions and mitigate hallucinations by providing accurate, environment-specific knowledge.
Working memory is integrated into agent nodes of ReAcTree through two mechanisms. First, the executable skill set $\mathcal{A}^{n}_{t}$ is augmented with special actions, recall location of <<movable object>>, enabling the agent to query the stored location of any movable object directly instead of actively exploring the environment. This recall action exclusively queries working memory. Second, working memory is automatically updated whenever an agent observes movable objects during interaction. For instance, if an agent opens a fridge and finds juice, it updates the working memory to reflect that juice is now in the fridge. We implement working memory as a lightweight Python dictionary that maps object classes to lists of observed instances and their associated locations (e.g., IDs, rooms, receptacles). This mechanism extends the concept of tool usage in language models (Schick et al., 2024), where the agent treats recall location as a specialized tool-like action.

### v2

@@ Abstract.
Recent advancements in large language models (LLMs) have enabled significant progress in decision-making and task planning for embodied autonomous agents. However, most existing methods struggle with complex, long-horizon tasks because they rely on a monolithic trajectory that entangles all past decisions and observations to solve the entire task in a single unified process. To address this limitation, we propose ReAcTree, a hierarchical task-planning method that decomposes a complex goal into manageable subgoals within a dynamically constructed agent tree. Each subgoal is handled by an LLM agent node capable of reasoning, acting, and further expanding the tree, while control flow nodes coordinate the execution strategies of agent nodes. In addition, we integrate two complementary memory systems: each agent node retrieves goal-specific, subgoal-level examples from episodic memory and shares environment-specific observations through working memory. Experiments on the WAH-NL and ALFRED show ReAcTree consistently outperforms strong task-planning baselines such as ReAct across diverse LLMs. Notably, on WAH-NL, ReAcTree achieves a 61% goal success rate with Qwen 2.5 72B, nearly doubling ReAct’s 31%. The code is available at https://github.com/Choi-JaeWoo/ReAcTree.git.
@@
Hierarchical Task Planning; LLM Agents; Behavior Trees
doi:
UCGT7089
1. Introduction
Large language models (LLMs) have recently exhibited remarkable reasoning capabilities by generating intermediate reasoning steps wei2022chain; kojima2022large. These advances open up new possibilities for decision-making and task planning in embodied autonomous agents, with the long-term objective of enabling them to fulfill high-level natural language commands autonomously. Early works huang2022language; ahn2022can; song2023llm; liang2023code have demonstrated that LLMs can leverage their pre-trained world knowledge to generate mid-level action sequences without additional training, primarily through prompting rather than handcrafted heuristics or learned policies eysenbach2019search; xu2019regression.
The decision-making and planning capabilities of LLMs can be further improved by incorporating observations or feedback from external environments. Recent approaches explore adapting the next action yao2023react; huang2022inner; zeng2023socratic, adjusting the entire plan shinn2024reflexion, or refining both the plan and the next action wang2023describe; sun2023adaplanner. However, these methods often use a monolithic trajectory that entangles all past decisions and observations from multiple subgoals, increasing the risk of hallucination and logical failures liu2024lost. Another line of research investigates multiple reasoning paths wang2023selfconsistency; yao2024tree; zhou2024language for general reasoning tasks. Such methods typically assume the capability to simulate multiple paths and revert to previous states zhao2024large; zhuang2024toolchain; hu2024treeplanner, which is rarely feasible in real-world scenarios due to irreversible actions (e.g., slicing an apple) and partial observability.
Figure 1. Illustration of how ReAcTree generates an agent tree for the instruction: Please bring one pudding and one juice to the coffee table. The left side shows the hierarchical structure with agent nodes (circles) and control flow nodes (squares); the number inside each circle denotes execution order, and the attached text box specifies the corresponding subgoal. Control flow nodes are labeled by type ($\rightarrow$ for sequence, ? for fallback, and $\Rightarrow$ for parallel). The right side highlights the decision-making trajectory of agent node 3, including observation, reasoning, acting, expansion, and episodic and working memory usage.
@@ 1. Introduction
The figure is divided into two parts illustrating the ReAcTree framework. The left side shows the hierarchical decomposition of a complex task into an agent tree using control flow nodes like sequence, fallback, and parallel. The right side zooms in on a single agent node to demonstrate its internal decision-making process, which involves retrieving examples from episodic memory, reasoning to generate actions, and dynamically expanding the tree structure.
@@
To tackle these challenges, we introduce ReAcTree, a novel framework that dynamically constructs an agent tree in the subgoal space rather than an action tree over primitive actions. Each LLM-powered agent node is responsible for a subgoal and extends the ReAct paradigm yao2023react: it can reason, act, or expand the tree by proposing new subgoals with an associated control flow when a task is too complex. Control flow nodes, inspired by Behavior Trees colledanchise2018learning, coordinate these agents by sequencing, fallback, or parallel execution. Conceptually, ReAcTree extends the Least-to-Most prompting strategy zhou2023least from static reasoning to dynamic, agentic planning. By decomposing a problem into semantically isolated subgoals and providing each with its own focused context, ReAcTree prevents the propagation of reasoning errors and makes long-horizon tasks more tractable for LLMs. The hierarchical structure of ReAcTree is illustrated in Figure 1.
@@
To enhance in-context learning and coordination capabilities, we incorporate two complementary memory systems. First, episodic memory assists each agent node in effective in-context learning by providing subgoal-level examples from past experiences that are semantically similar to the current subgoal. It can be bootstrapped from a small set of manually collected trajectories and gradually expanded through successful runs. Second, working memory functions as a shared blackboard, allowing agent nodes to exchange critical observations (e.g., the locations of movable objects). This collective situational awareness reduces redundant searches and mitigates hallucinations. Together, these memory systems enable ReAcTree to learn from the past and coordinate in the present, strengthening the robustness of agentic decision-making under partial observability.
@@
We evaluate ReAcTree by extending LoTa-Bench choi2024lotabench, a benchmark providing two household task environments (WAH-NL puig2021watchandhelp; choi2024lotabench with VirtualHome puig2018virtualhome and ALFRED shridhar2020alfred with AI2THOR kolve2017ai2), to a more challenging partially observable setting. This setup reflects real-world constraints and tests agentic decision-making under uncertainty. Across both environments, ReAcTree consistently outperforms strong baselines across various LLMs. On WAH-NL, it achieves a 61% goal success rate with Qwen 2.5 72B, nearly doubling ReAct’s 31% under the same model size. Notably, even with a much smaller 7B model, it reaches 37%, showing that ReAcTree maintains strong performance despite reduced model capacity. The contributions of its core components are further substantiated by ablation studies on memory systems and control flow types, as well as in-depth analyses of computational cost and failure modes.
@@
In summary, this paper makes the following contributions: (1) ReAcTree, a hierarchical planning framework for long-horizon tasks that dynamically constructs an LLM agent tree, where each agent node independently solves a subgoal while control flow nodes coordinate their execution into a coherent plan; (2) two memory systems: episodic memory that enhances in-context learning ability of agent nodes, and working memory that facilitates information sharing across agent nodes to support collective situational awareness; and (3) extensive experiments on LoTa-Bench under partially observable settings, demonstrating the effectiveness of ReAcTree across strong baselines and providing detailed analysis through ablations, computational cost analysis, and failure case studies.
@@
[…]
The figure diagrams the execution cycle in ReAcTree. It shows how an Agent Node chooses between reasoning, acting, or expanding into a subtree, and how the resulting Control Flow Node (Sequence, Fallback, or Parallel) manages the execution of child nodes.
@@ 2. Related Work
LLM-based Embodied Agents. LLMs such as GPT-3, GPT-4, PaLM, LLaMA, and Qwen brown2020language; achiam2023gpt; chowdhery2023palm; dubey2024llama; yang2024qwen2 have demonstrated impressive few-shot in-context learning capabilities, advancing the state of the art in NLP tasks. Prompting techniques wei2022chain; kojima2022large; zhou2023least; gao2023pal further enhance reasoning by encouraging articulation of intermediate steps. Building on these advancements, research has increasingly explored text-based applications to embodied decision-making. Early studies huang2022language; ahn2022can; song2023llm showed that LLMs could generate mid-level action sequences without additional training, while later studies introduced code-based plan generation liang2023code; singh2023progprompt, the extension of classical agent architectures ichida2024bdi, and the integration of environmental feedback or tools huang2022inner; zeng2023socratic; shen2024hugginggpt; lu2024chameleon; chen2023open. ReAct yao2023react enhanced planning by prompting explicit intermediate reasoning, while Reflexion shinn2024reflexion applied iterative self-refinement madaan2024self; welleck2023generating; paul2024refiner; chen2024teaching, offering additional flexibility for long-horizon tasks.
@@
Hierarchical Task Planning with LLMs. To tackle more complex, long-horizon tasks, researchers have introduced hierarchical frameworks that decompose goals into manageable planning levels. AdaPlanner sun2023adaplanner and DEPS wang2023describe adopt bi-level hierarchies, refining both an overall plan and next-step decisions with environmental feedback. Other methods combine classical task-and-motion planning with learned low-level controllers guided by LLM reasoning wong2024learning; liu2024learning; ding2023task; shukla2024lgts; wang2025inclet. Some works leverage structured formalisms, such as behavior trees wang2024mosaic; chen2024integrating; ao2025llm or hierarchical task networks munoz2025chathtn, to organize actions. However, they typically rely on predefined structures or domain-specific routines. In contrast, our approach enables dynamic task decomposition, expanding subgoals in complex environments without being tied to a specific domain.
@@
Tree Search-Based Planning with LLMs. A prominent line of research explores multiple reasoning paths to evaluate diverse hypotheses before committing to a final solution. Such approaches wang2023selfconsistency; yao2024tree; zhou2024language; xie2024self; besta2024graph; hao2023reasoning show that systematically branching reasoning steps significantly improves performance in reasoning tasks. Several studies extend this idea to agentic planning. For instance, LLM-MCTS zhao2024large and ToolChain* zhuang2024toolchain search action trees using Monte Carlo Tree Search and A* search, respectively. Tree-Planner hu2024treeplanner independently samples planning paths and merges them into an action tree, then makes decisions based on grounded observations at each node. However, these methods typically assume reversible actions (i.e., rolling back to prior states) in simulators, which limits their applicability in real-world settings. In contrast, ReAcTree constructs a dynamically expanding LLM agent tree that decomposes complex goals into subgoals, delegates them to agent nodes, and coordinates their execution via behavior tree-inspired control flow, replacing search-based exploration with agent coordination.
@@ 3. Preliminaries
Problem Formulation. We consider the agentic planning problem as a sequential decision-making problem aimed at achieving a goal $g$ expressed in natural language. At each time step $t$, the agent has access to the context $c_{t}=(o_{1},a_{1},o_{2},a_{2},\cdots,a_{t-1},o_{t})$, where $o_{i}$ and $a_{i}$ represent the observation and action at time step $i$, respectively. The objective of the agent is to generate the next action $a_{t}$ based on the context $c_{t}$, with the aim of eventually achieving the goal $g$.
ReAct (yao2023react). ReAct addresses this problem by interleaving reasoning and action execution using a pre-trained LLM $p_{LLM}$. The action policy is defined as $a_{t}\sim p_{LLM}(\cdot|P,g,c_{t})$, where $P=(P_{sys},P_{ic})$ is the initial prompt, composed of a system prompt $P_{sys}$ and in-context examples $P_{ic}$. The key idea is to use the augmented action space, $\hat{\mathcal{A}}_{t}=\mathcal{A}_{t}\cup\mathcal{L}$, where $\mathcal{A}_{t}$ is the set of executable skills available at time $t$, and $\mathcal{L}$ is the language space representing reasoning steps or thoughts. If $a_{t}\in\mathcal{A}_{t}$, the agent executes the action and obtains a text observation from the environment. If $a_{t}\in\mathcal{L}$, it is called a thought or reasoning trace, which aids in the logical inference of the LLM. In this case, the agent does not receive a new observation from the environment, i.e., $o_{t+1}=\phi$.
@@ 4. ReAcTree
ReAcTree is a hierarchical task-planning framework for agentic decision-making in complex environments. It dynamically constructs a tree of agent nodes and control flow nodes, as illustrated in Figure 2. Agent nodes operate as LLM-based task planners that can reason, act, and expand the tree with new subgoals and appropriate control flow. Control flow nodes, inspired by behavior trees (colledanchise2018learning), determine how child agent nodes are executed and how their outcomes are propagated upward in the tree. This design offers two key advantages: (1) isolating subgoals reduces hallucination and logical errors in long trajectories by keeping each agent node focused on its local context and goal, which also enables targeted in-context example selection, and (2) control flow nodes ensure robust and interpretable execution logic for reliable planning.
@@ 4.1. ReAcTree Algorithm
Control Flow Nodes. Each control flow node coordinates the execution of its child agent nodes according to behavior tree principles (colledanchise2018learning). ReAcTree supports three types of control flow nodes: sequence ($\rightarrow$), fallback (??), and parallel ($\Rightarrow$). The sequence node executes its children sequentially, returning success only if all of them succeed and failing immediately if any child fails. The fallback node also executes children in order but returns success as soon as the first child succeeds and fails only if all children fail. Lastly, the parallel node executes all children sequentially without early termination and aggregates outcomes according to a predefined policy; we adopt majority voting, useful for independent subgoals (e.g., placing multiple objects in different locations). After execution, a control flow node reports its overall success or failure to its parent node. The complete pseudocode is provided in Appendix A.
@@ 4.2. Memory Systems
To support ReAcTree’s hierarchical planning and decision-making, we introduce two complementary memory systems: episodic memory and working memory. Episodic memory enhances the in-context learning capability of agent nodes by storing and retrieving past subgoal-level experiences. In contrast, working memory facilitates information sharing across nodes by recording environment-specific observations, such as the latest location of movable objects.
@@
Episodic Memory. Each agent node $e$ handles a particular subgoal $g^{e}$, and its entire decision-making trajectory forms a subgoal-level experience. Episodic memory $M_{ep}$ stores experiences from agent nodes involved in task executions that ultimately succeed, even if the individual node did not complete its subgoal.
@@
By storing experiences at the subgoal level, episodic memory maintains shorter, goal-directed trajectories than monolithic planners like ReAct. For instance, while ReAct stores a single long trajectory for Bring pudding and juice to the coffee table, ReAcTree stores focused trajectories for each subgoal, such as find and pick up pudding and find and pick up juice. This granularity enables more targeted in-context retrieval and improves agent decision accuracy.
@@
Formally, each experience is recorded as a tuple $(t^{e},v^{e},s^{e})$, where: (1) $t^{e}=(g^{e},\,o^{e}_{1},\,a^{e}_{1},\,\dots,\,o^{e}_{T},\,a^{e}_{T})$ denotes the full text trajectory, recording all observations $o^{e}_{t}$ and actions $a^{e}_{t}$ during the node’s lifetime; (2) $v^{e}=f_{\text{sen}}(g^{e})$ is the sentence embedding of the node’s subgoal, computed using a pretrained encoder $f_{\text{sen}}$ such as Sentence-BERT (reimers2019sentence); and (3) $s^{e}\in\{\textit{success},\textit{failure},\textit{expand}\}$ is the termination state of the agent node.
@@
The collected episodic memory $M_{ep}$ provides in-context examples for agent node inference. Before decision-making begins at inference time, an agent node retrieves in-context examples from $M_{ep}$ by comparing its current subgoal $g^{n}$ to the stored goals using cosine similarity. Specifically, the agent embeds its goal $g^{n}$ as $v^{n}=f_{\text{sen}}(g^{n})$ and computes similarity with each $v^{e}$ in $M_{ep}$: $\text{sim}(v^{n},v^{e})=(v^{n}\cdot v^{e})/(\|v^{n}\|\cdot\|v^{e}\|)$. The agent retrieves the top $k$ similar experiences (with $k$ determined by a token limit) as in-context examples. If multiple experiences have the same similarity score, ReAcTree samples uniformly across termination states {success, failure, expand} to promote diversity. In practice, episodic memory can be bootstrapped from a few manually collected trajectories and gradually augmented with additional successful executions.
@@
Working Memory. Working memory captures environment-specific information within a single ReAcTree run, shared across all agent nodes to store and recall key observations. In this work, we focus on tracking movable object locations to reduce redundant interactions and mitigate hallucinations by providing accurate, environment-specific knowledge.
@@
Working memory is integrated into agent nodes of ReAcTree via two mechanisms. First, the executable skill set $\mathcal{A}^{n}_{t}$ is augmented with special actions, recall location of <<movable object>>, enabling agents to query the stored location of any movable object directly instead of actively exploring the environment. This recall action exclusively queries working memory. Second, working memory is automatically updated whenever an agent observes movable objects during interaction. For instance, if an agent opens a fridge and finds juice, it updates the working memory to reflect that juice is now in the fridge. We implement working memory as a lightweight Python dictionary that maps object classes to lists of observed instances and their associated locations (e.g., IDs, rooms, receptacles). This mechanism extends the concept of tool usage in language models (schick2024toolformer), treating recall location as a specialized tool-like action.